OpenStudy (anonymous):

What is the slope of the line perpendicular to the following line? x + 5y = 10

OpenStudy (anonymous):

First get the equation in slope intercept form: y=mx + b

OpenStudy (anonymous):

would you change it such as 5y=10-x

OpenStudy (anonymous):

Yes, now divide by the 5 to get y by itself.

OpenStudy (anonymous):

y=2-x

OpenStudy (anonymous):

You didn't divide the 5 for the x.